# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v3

@@ V The analysis of the continuous-time dynamics
To analyze the continuous-time dynamics, the equilibrium points of (7) should be characterized. The following subsection provides an intuition for how analyzing the evolution of $x(t)$ in Definition 6, i.e., the population proportion of strategy-$\mathtt{A}$ players, helps that of continuous-time population dynamics.
@@ V-A Intuition and example
${\mathcal{V}}_{1}(\bm{x})=\begin{cases}\{0.6-x_{1}\}&\text{if }x<0.85,\\ [-x_{1},0.6-x_{1}]&\text{if }x=0.85,\\ \{-x_{1}(t)\}&\text{if }x>0.85,\end{cases}$ (9a)
@@
  • The abstract dynamics. The thresholds divide the unit interval into four disjoint open intervals, i.e., $(0,0.35)$, $(0.35,0.75)$, $(0.75,0.85)$, and $(0.85,1)$. When $x\in(0,0.35)$, $\bm{\mathcal{V}}(\bm{x})$ is a singleton and $\dot{\bm{x}}$ will be equal to $(0.6-x_{1},-x_{2},-x_{3})$. The derivative of $x$, will then be $\dot{x}=0.6-x$. A similar procedure can be applied to the other intervals, resulting in $\dot{x}\in\mathcal{X}(x)$, where
@@
  • Abstract dynamics’ equilibria. At equilibrium, we have $0\in\mathcal{X}(x^{*})$. So we can find the equilibria by investigating all intervals in (10). The first two intervals are equilibrium-free. However, the image of the interval $(0.35,0.75)$ under $\mathcal{X}$ is $(-0.05,0.35)$, which includes zero. Therefore, the abstract dynamics admit an equilibrium point in this interval, which turns out to be $x^{*}_{1}=0.7$ as $0\in\mathcal{X}(0.7)$. Following the same procedure, we find two additional equilibrium points $x^{*}_{2}=0.75$ and $x^{*}_{3}=0.85$ for the abstract dynamics.
@@
Example 3 gives us an intuition that it is useful to focus on the evolution of the population proportion of $\mathtt{A}$-players, i.e., $x(t)$, which from now on we refer to as the abstract state as it is abstracting away the heterogeneity of the population. Define ${\eta}^{\prime}_{j}$ (resp. ${\eta}_{i}$) as the cumulative population proportion of coordinating (resp. anticoordinating) subpopulations, that is the population proportion of those having thresholds equal to or less (resp. greater) than $\tau^{\prime}_{j}$ (resp. $\tau_{i}$), i.e.,
@@
In words, it is assumed that the sum of the cumulative population proportion of first $k$ anticoordinating subpopulations and first $l$ coordinating subpopulations does not match any of the thresholds for the $\mathsf{p}+\mathsf{p}^{\prime}$ subpopulations. We refer the reader to Remark 4 for consequences of this assumption.
@@
[…]
where for $x=\tau_{i}$, $j$ equals $\max\{k|\tau^{\prime}_{k}<\tau_{i}\}$ and for $x=\tau_{j}^{\prime}$, $i$ equals $\max\{k|\tau^{\prime}_{j}<\tau_{k}\}$.
@@ V-C Equilibrium points
Inspired by the definitions in [44] and based on Example 3, we claim that the equilibria of the population dynamics are either clean-cut or ruffled as defined in the following.
@@
    ${\eta}_{i}+{\eta}^{\prime}_{j-1}<\tau^{\prime}_{j}<{\eta}_{i}+{\eta}^{\prime}_{j}.$ (14)
@@ V-D Global stability analysis
The abstract dynamics and, in turn, the continuous-time population dynamics may admit several clean-cut, anticoordinator-driven, or coordinator-driven equilibrium points. Denote the set of clean-cut, anticoordinator-driven, and coordinator-driven equilibrium points of the continuous-time population dynamics by $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$, respectively. When there is no equilibrium point of either type, the corresponding set will be empty. Assume that there are altogether $\mathtt{Q}\in\mathbb{N}$ equilibrium points of clean-cut or anticoordinator-driven type. Arrange the clean-cut or anticoordinator-driven equilibrium points of the continuous-time population dynamics (7) in the ascending order of their associated abstract equilibrium points. Hence, we have
@@
where $q^{*}_{k}$ is the abstract state at $\mathbf{q}^{*}_{k}\in\bm{\mathcal{Q}}^{a}\cup\bm{\mathcal{Q}}^{c}$ for $k\in[\mathtt{Q}]$. A set $\bm{\mathcal{M}}$ is attractive under differential inclusion (4) from a set $\bm{\mathcal{U}}$ if for each solution $\bm{x}(t)$ with $\bm{x}(0)\in\bm{\mathcal{U}}$ and each open $\epsilon$-neighborhood of $\bm{\mathcal{M}}$, there exists time $T>0$ such that $\bm{x}(t)$ falls in $\epsilon$-neighborhood of $\bm{\mathcal{M}}$ for all $t\geq T$ [45]. The union of all sets from which $\bm{\mathcal{M}}$ is attractive under differential inclusion $\bm{\mathcal{V}}$ is the basin of attraction of $\bm{\mathcal{M}}$ under $\bm{\mathcal{V}}$ [45].
According to A2, in the abstract dynamics, between every two anticoordinator-driven and/or clean-cut equilibria $q^{*}_{k}$ and $q^{*}_{k+1}$, $k\in[\mathtt{Q}-1]$, there exists exactly one coordinator-driven equilibrium, which we denote by $q^{*}_{k,k+1}$. The associated equilibrium in the continuous-time dynamics is denoted by $\mathbf{q}^{*}_{k,k+1}$. Define $q^{*}_{0,1}=0$ and $q^{*}_{\mathtt{Q},\mathtt{Q}+1}=1$. The following theorem reveals the global behavior of the continuous-time population dynamics.
@@
    The limit set of every point in $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x={q}^{*}_{k,k+1}\}$ for $k\in[\mathtt{Q}-1]$ is ${\{}\bm{q}^{*}_{k},\bm{q}^{*}_{k+1},\bm{q}^{*}_{k,k+1}{\}}$.
Indeed it can be shown that the clean-cut equilibrium points are exponentially asymptotically stable and the anticoordinator-driven equilibrium points are finite-time stable.
@@ V-E The Birkhoff center of the continuous-time population dynamics
The following corollaries are direct results of Proposition 3 and determine the Birkhoff center of the dynamical system induced by the continuous-time population dynamics capturing the behavior of a population of exclusively coordinators and exclusively anticoordinators, respectively.
@@
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics of a population of exclusively coordinators is
@@
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics describing the evolution of an exclusive population of $\mathsf{p}$ types of anticoordinators is
$\mathrm{BC}_{\text{A}}$ =$=\{\mathbf{q}\}.$ (17)
@@ VI The asymptotic behavior of the discrete population dynamics
1 ensures that $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$ are distinct. Otherwise, the following possible cases may happen: (i) The threshold of benchmark anticoordinating subpopulation $i$ equals ${\eta}^{\prime}_{j}+{\eta}_{i-1}$ (resp. ${\eta}^{\prime}_{j}+{\eta}_{i}$). In this case, following the approach taken in the proofs of Lemma 2 and Theorem 2, it can be shown that the equilibria $\mathbf{a}^{ij}$ and $\mathbf{c}^{(i-1)j}$ (resp. $\mathbf{c}^{ij}$) match and remain stable. (ii) The threshold of benchmark coordinating subpopulation $j$ equals ${\eta}^{\prime}_{j-1}+{\eta}_{i}$ (resp. ${\eta}^{\prime}_{j}+{\eta}_{i}$). In this case, following the approach taken in the proofs of Lemma 2 and Theorem 2, it can be shown that the equilibria $\mathbf{o}^{ij}$ and $\mathbf{c}^{i(j-1)}$ (resp. $\mathbf{c}^{ij}$) match, resulting in an unstable equilibrium. (iii) The threshold of benchmark coordinating subpopulation $j$ equals that of benchmark anticoordinating subpopulation $i$. In this case, the point $x=\tau_{i}$ is an equilibrium for the abstract dynamics if $\tau_{i}\in[{\eta}_{i-1}+{\eta}_{j-1}^{\prime},{\eta}_{i}+{\eta}_{j}^{\prime}]$. Following the approach taken in the proof of Theorem 2, it can be shown that if $\tau_{i}\in({\eta}_{i-1}+{\eta}_{j}^{\prime},{\eta}_{i}+{\eta}_{j-1}^{\prime})$, it is asymptotically stable, and unstable otherwise. In this case a continuum of equilibria $\mathbf{z}^{ij}$ emerges for the semicontinouous population dynamics. This continumm is characterized by $\bm{\mathcal{Z}}=\{\bm{x}\in\bm{\mathcal{X}}_{ss}|\forall l\notin\{i,\mathsf{p}+\mathsf{p}^{\prime}-j+1\},z^{ij}_{l}=a^{ij}_{l},z^{ij}_{i}+z^{ij}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}=\tau_{i}-({\eta}_{i-1}+{{\eta}_{j-1}^{\prime}})\}$. In words, all members of those coordinating subpopulations with thresholds less (resp. greater) than $\tau_{j}^{\prime}$ play strategy $\mathtt{A}$ (resp. $\mathtt{B}$), and all members of those anticoordinating subpopulations with thresholds less (resp. greater) than $\tau_{j}$ play strategy $\mathtt{B}$ (resp. $\mathtt{A}$). The proportions of strategy $\mathtt{A}$ players in the two benchmark subpopulations add up to $\tau_{i}-({\eta}_{i-1}+{{\eta}_{j-1}^{\prime}})$. The Birkhoff center will then also contain $\bm{\mathcal{Z}}$. The population proportion of strategy-$\mathtt{A}$ players is constant at $\bm{\mathcal{Z}}$ and consequently, the amplitude of the fluctuations in the population proportion of strategy-$\mathtt{A}$ players in the discrete population dynamics vanished with probability one.
Remark 5.
In some setups, such as in [15], agents include themselves when determining their preferred strategy. As detailed in Section VII-D, our main results hold for such update rules.
@@
Several studies have analyzed the population dynamics by studying one-dimensional dynamics. In [5], the discrete-time evolution of a heterogeneous population of coordinators was studied by analyzing the evolution of the population proportion of strategy-$\mathtt{A}$ players. A similar approach was used in [19, Section 9] for a heterogeneous population of coordinators and anticoordinators under a synchronous updating rule, where all agents are active at every time step. In [31, 32], scalar continuous-time dynamics instead serve as approximations of the evolution of the population proportion of strategy-$\mathtt{A}$ players in structured populations. In this paper, however, we were interested in studying the evolution of the proportion of strategy-$\mathtt{A}$ players within each subpopulation. Hence, a one-dimensional dynamic would not reveal the behavior of each sub-population. The dimension should be at least as many as the number of sub-populations, i.e., $\mathsf{p}+\mathsf{p}^{\prime}$. Moreover, the evolution of the total proportion of $\mathtt{A}$-players, i.e., $x^{\mathsf{N}}$, is not a Markov chain because the change from $x^{\mathsf{N}}$ to $x^{\mathsf{N}}+1/\mathsf{N}$ (or to $x^{\mathsf{N}}-1/\mathsf{N}$) depends on the proportions of $\mathtt{A}$-players in subpopulations. Nevertheless, the expected change of $x^{\mathsf{N}}$ can be calculated, and that is how we provided the abstract dynamics, which facilitated the subsequent analysis of the $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional population dynamics.
@@
Figure 4: Example 4. The solution of the continuous-time population dynamics for two different initial conditions. The upper panel shows the evolution of the population state which converges to clean-cut equilibrium point $\mathbf{c}^{22}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,3,0},{0,0,3,3})/28$ and the lower one depicts the evolution of the population state converging to the anticoordinator-driven equilibrium point $\mathbf{a}^{24}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,2,0},{6,8,3,3})/28$, where the corresponding population proportion of $\mathtt{A}$-players equals $\tau_{2}=6/7$. The observed non-smoothness in the evolution of some subpopulations is due to the change in their preferred strategy. For example, in the upper panel, at time zero, the abstract state is below $\tau^{\prime}_{2}$ (resp. $\tau_{3}$) resulting in $\dot{x}^{\prime}_{2}<0$ ($\dot{x}_{3}>0$). As abstract state increases and exceeds $\tau^{\prime}_{2}$ (resp. $\tau_{3}$), the preferred strategy of coordinating (resp. anticoordinating) subpopulation $2$ (resp. $3$) changes to $\mathtt{A}$ (resp. $\mathtt{B}$) yielding $\dot{x}^{\prime}_{2}>0$ (resp. $\dot{x}_{3}<0$).
@@ VII Concluding Remarks
The convergence result of the continuous-time population dynamics for exclusive populations of anticoordinators, i.e., 3, matches that of the discrete dynamics [16] where the population state was proven to always reach the clean-cut equilibrium $\mathbf{q}$ or reach the set of two ruffled states (that are not possible here due to the continuity of the state space of the continuous-time population dynamics). Similarly, the convergence result of the continuous-time population dynamics for exclusive populations of coordinators, i.e., 2, matches that of the discrete dynamics [15] where the population state was proven to always equilibrate. Moreover, the equilibria of the discrete and continuous-time dynamics match.
@@
In addition, the stability of clean-cut equilibrium points of continuous-time population dynamics matches that of the discrete dynamics [46]. However, the stability of the ruffled types has not been investigated in the discrete population.
@@
In case of the fluctuations, in all of the examples, we observed that the interval defined by the minimum and maximum of the ratio of strategy $\mathtt{A}$-players in the same population and for a fixed initial condition and different activation sequence contains an anticoordinator-driven equilibrium point of the abstract dynamics. Moreover, as the population size grows, the interval shrinks. It remains yet unknown whether this observation holds true always for every population and whether ruffled anticoordinator-driven equilibria are necessary for the existence of minimal positively invariant sets in the finite discrete populations. The difficulty partly arises from the inherent challenge of characterizing the positively invariant sets and amplitude of the fluctuations. Although [19, Corollary 6] provides bounds on the fluctuations, it still requires the benchmark subpopulations to be known, which appears to be an unsolved complicated problem. The result of this paper holds true in a well-mixed population. Whether a similar result can be obtained for structured populations, such as k-regular networks, is an open question. In this regard, approaches similar to those in [32] might work.

### v4

@@ V The analysis of the continuous-time dynamics
To analyze the continuous-time dynamics, the equilibrium points of (7) should be characterized. The following subsection provides an intuition for how analyzing the evolution of $x(t)$ in Definition 6, i.e., the population proportion of $\mathtt{A}$-players, helps that of continuous-time population dynamics.
@@ V-A Intuition and example
${\mathcal{V}}_{1}(\bm{x})=\begin{cases}\{0.6-x_{1}\}&\text{if }x<0.85,\\ [-x_{1},0.6-x_{1}]&\text{if }x=0.85,\\ \{-x_{1}\}&\text{if }x>0.85,\end{cases}$ (9a)
@@
  • The abstract dynamics. The thresholds divide the unit interval into four disjoint open intervals, i.e., $(0,0.35)$, $(0.35,0.75)$, $(0.75,0.85)$, and $(0.85,1)$. When $x\in(0,0.35)$, $\bm{\mathcal{V}}(\bm{x})$ is a singleton and $\dot{\bm{x}}$ will be equal to $(0.6-x_{1},-x_{2},-x_{3})$. The derivative of $x$ will then be $\dot{x}=0.6-x$. A similar procedure can be applied to the other intervals, resulting in $\dot{x}\in\mathcal{X}(x)$, where
@@
  • Abstract dynamics’ equilibria. At equilibrium, we have $0\in\mathcal{X}(x^{*})$. So we can find the equilibria by investigating all intervals in (10). The first two intervals are equilibrium-free. However, the image of the interval $(0.35,0.75)$ under $\mathcal{X}$ is (0.05,0.35)(-0.05,0.35), which includes zero. Therefore, the abstract dynamics admit an equilibrium point in this interval, which turns out to be $x^{*}_{1}=0.7$ as $0\in\mathcal{X}(0.7)$. Following the same procedure, we find two additional equilibrium points $x^{*}_{2}=0.75$ and $x^{*}_{3}=0.85$ for the abstract dynamics.
@@
Example 3 gives us an intuition that it is useful to focus on the evolution of the population proportion of $\mathtt{A}$-players, i.e., $x(t)$, which from now on we refer to as the abstract state as it is abstracting away the heterogeneity of the population. Define ${\eta}^{\prime}_{j}$ (resp. ${\eta}_{i}$) as the cumulative population proportion of coordinating (resp. anticoordinating) subpopulations, that is, the population proportion of those having thresholds equal to or less (resp. greater) than $\tau^{\prime}_{j}$ (resp. $\tau_{i}$), i.e.,
@@
In words, it is assumed that the sum of the cumulative population proportion of the first $k$ anticoordinating subpopulations and first $l$ coordinating subpopulations does not match any of the thresholds for the $(\mathsf{p}+\mathsf{p}^{\prime})$ subpopulations.
@@
$\mathcal{X}(x)=\begin{cases}[{\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]\hskip 50.0pt\text{if }\exists i(x=\tau_{i}),&\\ [{\eta}_{i}+{\eta}^{\prime}_{j-1}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]\hskip 50.0pt\text{if }\exists j(x=\tau^{\prime}_{j}),&\\ \{{\eta}_{i}+{\eta}^{\prime}_{j}-x\}\\ \quad\qquad\text{if }\exists i,j(x\in(\max\{\tau^{\prime}_{j},\tau_{i+1}\},\min\{\tau^{\prime}_{j+1},\tau_{i}\})),\end{cases}$
where for $x=\tau_{i}$, $j=\max\{k\in[\mathsf{p}^{\prime}]\cup\{0\}|\tau^{\prime}_{k}<\tau_{i}\}$ and for $x=\tau_{j}^{\prime}$, $i$ is equal to $\max\{k\in[\mathsf{p}]\cup\{0\}|\tau^{\prime}_{j}<\tau_{k}\}$.
@@ V-C Equilibrium points
Inspired by the definitions in [ramazi2017asynchronous] and based on Example 3, we claim that the equilibria of the population dynamics are either clean-cut or ruffled as defined in the following.
@@
    $0<\tau^{\prime}_{j}-({\eta}_{i}+{\eta}^{\prime}_{j-1})<{\rho}^{\prime}_{j}.$ (14)
@@ V-D Global stability analysis
The abstract dynamics and, in turn, the continuous-time population dynamics may admit several clean-cut, anticoordinator-driven, or coordinator-driven equilibrium points.
Denote the set of clean-cut, anticoordinator-driven, and coordinator-driven equilibrium points of the continuous-time population dynamics by $\bm{\mathcal{Q}}^{c}$, $\bm{\mathcal{Q}}^{a}$, and $\bm{\mathcal{Q}}^{o}$, respectively. When there is no equilibrium point of either type, the corresponding set will be empty. Assume that there are altogether $\mathtt{Q}\in\mathbb{N}$ equilibrium points of clean-cut or anticoordinator-driven type. Arrange the clean-cut or anticoordinator-driven equilibrium points of the continuous-time population dynamics (7) in the ascending order of their associated abstract equilibrium points. Hence, we have
@@
[…]
According to A2, between every two consecutive anticoordinator-driven and/or clean-cut equilibria $q^{*}_{k}$ and $q^{*}_{k+1}$, $k\in[\mathtt{Q}-1]$, there exists exactly one coordinator-driven equilibrium, which we denote by $q^{*}_{k,k+1}$. The associated equilibrium in the continuous-time dynamics is denoted by $\mathbf{q}^{*}_{k,k+1}$. Define $q^{*}_{0,1}=0$ and $q^{*}_{\mathtt{Q},\mathtt{Q}+1}=1$. The following theorem reveals the global behavior of the continuous-time population dynamics.
@@
    The limit set of every point in the set $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x={q}^{*}_{k,k+1}\}$ for $k\in[\mathtt{Q}-1]$ is ${\{}\bm{q}^{*}_{k},\bm{q}^{*}_{k+1},\bm{q}^{*}_{k,k+1}{\}}$.
Indeed, it can be shown that the clean-cut equilibrium points are exponentially asymptotically stable and the anticoordinator-driven equilibrium points are finite-time stable.
@@ V-E The Birkhoff center of the continuous-time population dynamics
The following corollaries are direct results of the Proposition 3 and determine the Birkhoff centers of the dynamical systems induced by the continuous-time population dynamics capturing the behavior of the populations of all coordinators and all anticoordinators, respectively.
@@
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics associated with a population of all coordinators is
@@
If Assumption 1 holds, the Birkhoff center of the continuous-time population dynamics describing the evolution of a population of $\mathsf{p}$ types of anticoordinators is
$\mathrm{BC}_{\text{A}}$ $=\{\mathbf{q}\}.$ (17)
@@ VI The asymptotic behavior of the discrete population dynamics
In some setups, such as in [ramazi2020convergence], agents include themselves when determining their preferred strategy. Our main results hold for such update rules.
@@
Several studies have analyzed the population dynamics by studying one-dimensional dynamics. In [granovetter1978threshold], the discrete-time evolution of a heterogeneous population of coordinators was studied by analyzing the evolution of the population proportion of strategy-$\mathtt{A}$ players. A similar approach was used in [roohi, Section 9] for a heterogeneous population of coordinators and anticoordinators under a synchronous updating rule, where all agents are active at every time step. In [comoApproximation, ravazzi2023asynchronous], scalar continuous-time dynamics instead serve as approximations of the evolution of the population proportion of strategy-$\mathtt{A}$ players in structured populations. In this paper, however, we were interested in studying the evolution of the proportion of strategy-$\mathtt{A}$ players within each subpopulation. Hence, a one-dimensional dynamic would not reveal the behavior of each sub-population. The dimension should be at least as many as the number of sub-populations, i.e., $\mathsf{p}+\mathsf{p}^{\prime}$. Moreover, the evolution of the total proportion of $\mathtt{A}$-players at index $k$ depends on the proportions of $\mathtt{A}$-players in subpopulations. Nevertheless, the expected change of $x^{\mathsf{N}}$ can be calculated, and that is how we provided the abstract dynamics, which facilitated the subsequent analysis of the $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional population dynamics.
@@
Figure 4: Example 4. The solution of the continuous-time population dynamics for two different initial conditions. The upper panel shows the evolution of the population state which converges to clean-cut equilibrium point $\mathbf{c}^{22}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,3,0},{0,0,3,3})/28$ and the lower one depicts the evolution of the population state converging to the anticoordinator-driven equilibrium point $\mathbf{a}^{24}=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2,2,0},{6,8,3,3})/28$, where the corresponding population proportion of $\mathtt{A}$-players equals $\tau_{2}=6/7$. The observed non-smoothness in the evolution of some subpopulations is due to the change in their preferred strategy. For example, in the upper panel, at time zero, the abstract state is below $\tau^{\prime}_{2}$ (resp. $\tau_{3}$), resulting in $\dot{x}^{\prime}_{2}<0$ ($\dot{x}_{3}>0$). As abstract state increases and exceeds $\tau^{\prime}_{2}$ (resp. $\tau_{3}$), the preferred strategy of coordinating (resp. anticoordinating) subpopulation $2$ (resp. $3$) changes to $\mathtt{A}$ (resp. $\mathtt{B}$) yielding $\dot{x}^{\prime}_{2}>0$ (resp. $\dot{x}_{3}<0$).
@@ VII Concluding Remarks
The convergence result of the continuous-time population dynamics for populations of all anticoordinators, i.e., 3, matches that of the discrete dynamics [ramazi2017asynchronous] where the population state was proven to always reach the clean-cut equilibrium $\mathbf{q}$ or reach the set of two ruffled states (that are not possible here due to the continuity of the state space of the continuous-time population dynamics). Similarly, the convergence result of the continuous-time population dynamics for populations of all coordinators, i.e., 2, matches that of the discrete dynamics [ramazi2020convergence] where the population state was proven to always equilibrate. Moreover, the equilibria of the discrete and continuous-time dynamics match.
@@
In addition, the stability of clean-cut equilibrium points of continuous-time population dynamics matches that of the discrete dynamics [le2022heterogeneous]. However, the stability of the ruffled types has not been investigated in the discrete population.
@@
In the case of the fluctuations, in all examples, we observed that the interval defined by the minimum and maximum of the ratio of strategy $\mathtt{A}$-players in the same population and for a fixed initial condition and different activation sequence contains an anticoordinator-driven equilibrium point of the abstract dynamics. Moreover, as the population size grows, the interval shrinks. It remains unknown whether this observation holds always for every population and whether ruffled anticoordinator-driven equilibria are necessary for the existence of minimal positively invariant sets in the finite discrete populations. The difficulty partly arises from the inherent challenge of characterizing the positively invariant sets and amplitude of the fluctuations. Although [roohi, Corollary 6] provides bounds on the fluctuations, it still requires the benchmark subpopulations to be known, which appears to be an unsolved complicated problem. The result of this paper holds in a well-mixed population. Whether a similar result can be obtained for structured populations, such as k-regular networks, is an open question. In this regard, approaches similar to those in [ravazzi2023asynchronous] might work.
Acknowledgment
The authors thank the anonymous reviewers for constructive comments and Hichem Ben-El-Mechaiekh for technical discussions.